Genetic Insights Co. purchases an asset for $11,585. This asset qualifies as a seven-year recovery asset under MACRS. The seven-year fixed depreciation percentages for years 1, 2, 3, 4, 5, and 6 are 14.29%, 24.49%, 17.49%, 12.49%, 8.93%, and 8.93%, respectively. Genetic Insights has a tax rate of 30%. The asset is sold at the end of six years for $3,594. Calculate After-Tax Cash Flow at disposal.

Round the answer to two decimals.
Answer #1

Book value as on date of sale=Cost-Accumulated Depreciation

=11,585*(1-0.1429-0.2449-0.1749-0.1249-0.0893-0.0893)

=$1550.073

Hence gain on sale=3594-1550.073=$2043.927

Hence after-tax cash flow at disposal=Sale proceeds-(Tax rate*Gain on sale)

=3594-(2043.927*0.3)

=$2980.82(Approx).